[figure: 9.7 cm]

HYPERFINE PARAMETERS OF THE HEUSLER TYPE Fe_{2+x}V_{1−x} ALLOYS

J. Deniszczyk a Institute of Physics and Chemistry of Metals, University of Silesia 40-007 Katowice, Bankowa 12, Poland

The hyperfine parameters (isomer shift and hyperfine magnetic field) for ^{57}Fe nucleus in the Heulser-type Fe_{2+x}V_{1−x} (x ≥ 0) alloys were calculated by means of the ab initio electronic structure calculation methods. The calculations performed for these alloys have shown that the Fe anti-site (AS) atoms (Fe at V positions of the Heulser Fe_{2}VAl) characterizes the stable (independent on concentration x) magnetic moment of magnitude of 2.2—2.5 μ_{B}. On the other hand the Mössbauer experimental investigations of the Fe_{2+x}V_{1−x} (x ≥ 0) alloys have indicated the presence of Fe atoms with intermediate value of hyperfine magnetic field (∼15 T). The experimental data interpreted within the simple model of linear scaling (HF∼ μ_{F e}) have predicted the presence of the Fe atoms with the medium sized magnetic moment (∼1.5 μ_{B}) which number varies with temperature and concentration. The results of calculations for the hyperfine magnetic field (HF) have explained the experimental data in the following way. The calculated magnitude of HF on Fe-AS atoms depends sensitively on the magnetization of the nearest neighborhs. The transferred part of the HF, usually small as compared the local part, reaches relatively large values and points toward opposite direction compared the local HF. The both parts of HF (local and transferred) reduce each other on Fe-AS atoms leading to the medium sized magnitude of HF despite the high value of magnetic moment of these atoms.

[figure: 13.4 cm]

Subject category :

3. Transition Metals, Alloys and Compounds

Presentation mode : poster

Corresponding author : J. Deniszczyk

Address for correspondence : Institute of Physics and Chemistry of Metals University of Silesia 40-007 Katowice, Bankowa 12

Email address : jdeni@us.edu.pl